if a missle is launched at a 30 degree angle and lands 7500 km away what is the formula to find the area under that arc
Answer #1

Area under the arc = pi*r^2 / 2

where r = 1/2 * (distance the missile lands)

= pi*(7500/2 * 1000) / 2

= 2.2*10^13 m2

so the area under the arc = 2.2*10^13 m^2